# Two-Timescale Gradient Descent Ascent Algorithms for Nonconvex Minimax Optimization

### v2

@@ 3 Preliminaries and Technical Background
The stationarity of $f$ is equivalent to Definition 3.10 up to a constant. Indeed, if $(\mathbf{x},\mathbf{y})\in\mathbb{R}^{m}\times\mathcal{Y}$ is an $\epsilon$-stationary point as per Definition 3.10, we have $\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\|\leq\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y}^{+})\|+\ell\|\mathbf{y}^{+}-\mathbf{y}\|\leq 2\epsilon$. Conversely, if $(\mathbf{x},\mathbf{y})\in\mathbb{R}^{m}\times\mathcal{Y}$ satisfies $\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\|\leq\epsilon$ and $\|\mathbf{y}^{+}-\mathbf{y}\|\leq\frac{\epsilon}{\ell}$, we have $\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y}^{+})\|\leq\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\|+\ell\|\mathbf{y}^{+}-\mathbf{y}\|\leq 2\epsilon$.
@@
Propositions 3.11 and 3.12 show that one needs to pay an additional $O(\kappa\log(1/\epsilon))$ or $O(\epsilon^{-2})$ gradient evaluations for translating the stationarity of $f$ to that of $\Phi$ in nonconvex-strongly-concave and nonconvex-concave settings. In this sense, the notions of stationarity as per Definitions 3.3 and 3.8 are stronger than the notion of stationarity as per Definition 3.10.
@@ 4.1 Main results
    in the TTGDA scheme (cf. Algorithm 1), the required number of gradient evaluations to return an $\epsilon$-stationary point is bounded by
@@
    If we set $\eta_{\mathbf{x}}^{t}$ and $\eta_{\mathbf{y}}^{t}$ as per Eq. (4.1), $M=\max\{1,48\kappa\sigma^{2}\epsilon^{-2}\}$ and $G$ as per Eq. (1.3) in the TTSGDA scheme (cf. Algorithm 2), the required number of stochastic gradient evaluations to return an $\epsilon$-stationary point is bounded by
@@
    in the TTGDA scheme (cf. Algorithm 1), the required number of gradient evaluations to return an $\epsilon$-stationary point is bounded by
@@
    in the TTSGDA scheme (cf. Algorithm 2), the required number of stochastic gradient evaluations to return an $\epsilon$-stationary point is bounded by
@@ 5.1 Main results
    in the TTGDA scheme (cf. Algorithm 1), the required number of gradient evaluations to return an $\epsilon$-stationary point is bounded by
@@
    in the TTSGDA scheme (cf. Algorithm 2), the required number of stochastic gradient evaluations to return an $\epsilon$-stationary point is bounded by
@@
    in the TTGDA scheme (cf. Algorithm 1), the required number of gradient evaluations to return an $\epsilon$-stationary point is bounded by
@@
    in the TTSGDA scheme (cf. Algorithm 2), the required number of stochastic gradient evaluations to return an $\epsilon$-stationary point is bounded by
@@ 6.3 WGANs with nonlinear generators
Figure 3 shows that TTSGDA is worse with Adam and RMSprop in this complex setting. This makes sense since the nonlinear generators result in a richer structure and the adaptive algorithms are highly optimized for handling such structure. The vanilla implementation of TTSGDA does not achieve a satisfactory level of performance and can be enhanced by exploiting the advanced tools of regularization and adaptiveness (Yang et al., 2022a, b; Li et al., 2023). For more details on the implementation heuristics and empirical performance for TTSGDA in training GANs and its variants, we refer to Heusel et al. (2017) for the details.
@@ Appendix A Proofs for Technical Lemmas
To further facilitate the reader’s understanding, we give a brief description of the extragradient algorithm. It was introduced by Korpelevich [1976] and has been recognized as one of the most useful methods for solving minimax optimization problems. In the convex-concave setting, its global convergence rate guarantees were established for both averaged iterates [Nemirovski, 2004, Mokhtari et al., 2020a, b] and last iterates [Gorbunov et al., 2022, Cai et al., 2022]. When applied to solve the optimization problem of $\min_{\mathbf{x}}\max_{\mathbf{y}\in\mathcal{Y}}f(\mathbf{x},\mathbf{y})+\ell\|\mathbf{x}-\hat{\mathbf{x}}\|^{2}$, one iteration of the extragradient algorithm is summarized as follows,
@@
By applying the same argument as used for proving Lemma 4.3, we have that $\partial_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\subseteq\partial\Phi(\mathbf{x})$ for each $\mathbf{x}$ where $\mathbf{y}\in\mathcal{Y}^{\star}(\mathbf{x})$.
@@ Appendix B Smooth and Nonconvex-Strongly-Concave Setting
We now prove the results for Algorithm 2. By applying the same argument as used for proving Eq. (B.1), we have
@@
This implies that the number of gradient evaluations required by Algorithm 1 to return an $\epsilon$-stationary point is bounded by
@@
We then prove the results for Algorithm 2. By applying the same argument as used for analyzing Algorithm 1 but with the second inequalities in Lemmas B.2 and B.3, we have
@@
This implies that the number of stochastic gradient evaluations required by Algorithm 2 to return an $\epsilon$-stationary point is bounded by
@@ Appendix C Smooth and Nonconvex-Concave Setting
By applying the same argument as used for proving Eq. (C.1), we have
@@ A $=$ $\mathbb{E}[f(\mathbf{x}_{t-1},\mathbf{y}_{t})-f(\mathbf{x}_{t},\mathbf{y}_{t})],$
By applying same argument as used for analyzing Algorithm 1, we have
@@
We then prove the results for Algorithm 2. By applying the same argument as used for analyzing Algorithm 1 but with the second inequality in Lemma C.2, we obtain the inequality for Algorithm 2. $\Box$
@@
This implies that the number of gradient evaluations required by Algorithm 1 to return an $\epsilon$-stationary point is bounded by
@@
We then prove the results for Algorithm 2. By applying the same argument as used for analyzing Algorithm 1 but with the second inequalities in Lemmas C.1 and C.3, we have
@@
This implies that the number of stochastic gradient evaluations required by Algorithm 2 to return an $\epsilon$-stationary point is bounded by
@@ Appendix D Nonsmooth and Nonconvex-Strongly-Concave Setting
Proof. We first prove the results for Algorithm 1. By applying the same argument as used for proving Lemma C.1, we have
@@
Again, by applying the same argument as used for proving Lemma C.1, we have
@@
We then prove the results for Algorithm 2. By applying the same argument as used for proving Lemma C.1, we have
@@
Since $f(\cdot,\mathbf{y})$ is $L$-Lipschitz for each $\mathbf{y}\in\mathcal{Y}$, we have $\|\mathbf{g}\|\leq L$ for all $\mathbf{g}\in\partial_{\mathbf{x}}f(\mathbf{x},\mathbf{y})$ and each pair of $\mathbf{x}$ and $\mathbf{y}\in\mathcal{Y}$. Applying the same argument as used for proving Lemma C.2 yields
@@
Applying the same argument as used for proving Lemma C.2 yields
@@
Proof. By applying the same argument as used for proving Lemma C.3 but with the two inequalities in Lemma D.2, we can obtain the inequalities for Algorithm 1 and 2. $\Box$
@@
We first prove the results for Algorithm 1. By applying the same argument as used for proving Theorem 4.6 but with the first inequalities in Lemmas D.1 and D.3, we have
@@
This implies that the number of gradient evaluations required by Algorithm 1 to return an $\epsilon$-stationary point is bounded by
@@
We then prove the results for Algorithm 2. By applying the same argument as used for analyzing Algorithm 1 but with the second inequalities in Lemmas D.1 and D.3 and using $B=\lfloor\sqrt{\frac{1}{\mu\eta_{\mathbf{x}}}}\rfloor+1$ and the definition of $\eta_{\mathbf{x}}$, we have
@@
This implies that the number of stochastic gradient evaluations required by Algorithm 2 to return an $\epsilon$-stationary point is bounded by
@@ Appendix E Nonsmooth and Nonconvex-Concave Setting
Proof. By applying the same argument as used for proving Lemma C.3 but with the two inequalities in Lemma E.1, we can obtain the inequalities for Algorithm 1 and 2. $\Box$
@@
We first prove the results for Algorithm 1. By applying the same argument as used for proving Theorem 4.6 but with the first inequalities in Lemmas D.1 and E.2, we have
@@
This implies that the number of gradient evaluations required by Algorithm 1 to return an $\epsilon$-stationary point is bounded by
@@
We now prove the results for Algorithm 2. By applying the same argument as used for analyzing Algorithm 1 but with the second inequalities in Lemmas D.1 and E.2, we have
@@
This implies that the number of stochastic gradient evaluations required by Algorithm 2 to return an $\epsilon$-stationary point is bounded by

### v3

@@ 3 Preliminaries and Technical Background
The stationarity of $f$ is in fact equivalent to Definition 3.10 up to a constant. Indeed, if $(\mathbf{x},\mathbf{y})\in\mathbb{R}^{m}\times\mathcal{Y}$ is an $\epsilon$-stationary point as per Definition 3.10, we have $\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\|\leq\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y}^{+})\|+\ell\|\mathbf{y}^{+}-\mathbf{y}\|\leq 2\epsilon$. Conversely, if $(\mathbf{x},\mathbf{y})\in\mathbb{R}^{m}\times\mathcal{Y}$ satisfies $\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\|\leq\epsilon$ and $\|\mathbf{y}^{+}-\mathbf{y}\|\leq\frac{\epsilon}{\ell}$, we have $\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y}^{+})\|\leq\|\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\|+\ell\|\mathbf{y}^{+}-\mathbf{y}\|\leq 2\epsilon$.
@@
Propositions 3.11 and 3.12 show that one needs to pay an additional $O(\kappa\log(1/\epsilon))$ or $O(\epsilon^{-2})$ gradient evaluations for translating the stationarity of $f$ to that of $\Phi$ in nonconvex-strongly-concave and nonconvex-concave settings. In this sense, the notions of stationarity as per Definitions 3.3 and 3.8 are stronger than that as per Definition 3.10.
@@ 4.1 Main results
    in the TTGDA scheme (see Algorithm 1), the required number of gradient evaluations to return an $\epsilon$-stationary point is
@@
    If we set $\eta_{\mathbf{x}}^{t}$ and $\eta_{\mathbf{y}}^{t}$ as per Eq. (4.1), $M=\max\{1,48\kappa\sigma^{2}\epsilon^{-2}\}$ and $G$ as per Eq. (1.3) in the TTSGDA scheme (see Algorithm 2), the required number of stochastic gradient evaluations to return an $\epsilon$-stationary point is
@@
    in the TTGDA scheme (see Algorithm 1), the required number of gradient evaluations to return an $\epsilon$-stationary point is
@@
    in the TTSGDA scheme (see Algorithm 2), the required number of stochastic gradient evaluations to return an $\epsilon$-stationary point is
@@ 5.1 Main results
    in the TTGDA scheme (see Algorithm 1), the required number of gradient evaluations to return an $\epsilon$-stationary point is
@@
    in the TTSGDA scheme (see Algorithm 2), the required number of stochastic gradient evaluations to return an $\epsilon$-stationary point is
@@
    in the TTGDA scheme (see Algorithm 1), the required number of gradient evaluations to return an $\epsilon$-stationary point is
@@
    in the TTSGDA scheme (see Algorithm 2), the required number of stochastic gradient evaluations to return an $\epsilon$-stationary point is
@@ 6.3 WGANs with nonlinear generators
Figure 3 shows that TTSGDA is worse with Adam and RMSprop in this complex setting. This makes sense since the nonlinear generators result in a richer structure and the adaptive algorithms are highly optimized for handling such structure. The vanilla implementation of TTSGDA does not achieve a satisfactory level of performance and can be enhanced by exploiting the advanced tools of regularization and adaptiveness (Yang et al., 2022a, b; Li et al., 2023). For more details on the implementation heuristics and empirical performance for TTSGDA in training GANs and its variants, we refer to Heusel et al. (2017).
@@ Appendix A Proofs for Technical Lemmas
To facilitate the reader’s understanding, we give a brief description of the extragradient algorithm. It was introduced by Korpelevich [1976] and has been recognized as one of the most useful methods for solving minimax optimization problems. In the convex-concave setting, its global convergence rate guarantees were established for both averaged iterates [Nemirovski, 2004, Mokhtari et al., 2020a, b] and last iterates [Gorbunov et al., 2022, Cai et al., 2022]. When applied to solve the optimization problem of $\min_{\mathbf{x}}\max_{\mathbf{y}\in\mathcal{Y}}f(\mathbf{x},\mathbf{y})+\ell\|\mathbf{x}-\hat{\mathbf{x}}\|^{2}$, one iteration of the extragradient algorithm is summarized as follows,
@@
Applying the same argument used for proving Lemma 4.3, we have that $\partial_{\mathbf{x}}f(\mathbf{x},\mathbf{y})\subseteq\partial\Phi(\mathbf{x})$ for each $\mathbf{x}$ where $\mathbf{y}\in\mathcal{Y}^{\star}(\mathbf{x})$.
@@ Appendix B Smooth and Nonconvex-Strongly-Concave Setting
We now prove the results for Algorithm 2. Applying the same argument used for proving Eq. (B.1), we have
@@
This implies that the number of gradient evaluations required by Algorithm 1 to return an $\epsilon$-stationary point is
@@
We then prove the results for Algorithm 2. Applying the same argument used for analyzing Algorithm 1 but with the second inequalities in Lemmas B.2 and B.3, we have
@@
This implies that the number of stochastic gradient evaluations required by Algorithm 2 to return an $\epsilon$-stationary point is
@@ Appendix C Smooth and Nonconvex-Concave Setting
Applying the same argument used for proving Eq. (C.1), we have
@@ A $=$ $\mathbb{E}[f(\mathbf{x}_{t-1},\mathbf{y}_{t})-f(\mathbf{x}_{t},\mathbf{y}_{t})],$
Applying the same argument used for analyzing Algorithm 1, we have
@@
We then prove the results for Algorithm 2. Applying the same argument used for analyzing Algorithm 1 but with the second inequality in Lemma C.2, we can obtain the inequality for Algorithm 2. $\Box$
@@
This implies that the number of gradient evaluations required by Algorithm 1 to return an $\epsilon$-stationary point is
@@
We then prove the results for Algorithm 2. Applying the same argument used for analyzing Algorithm 1 but with the second inequalities in Lemmas C.1 and C.3, we have
@@
This implies that the number of stochastic gradient evaluations required by Algorithm 2 to return an $\epsilon$-stationary point is
@@ Appendix D Nonsmooth and Nonconvex-Strongly-Concave Setting
Proof. We first prove the results for Algorithm 1. Applying the same argument used for proving Lemma C.1, we have
@@
Applying the same argument used for proving Lemma C.1, we have
@@
We then prove the results for Algorithm 2. Applying the same argument used for proving Lemma C.1, we have
@@
Since $f(\cdot,\mathbf{y})$ is $L$-Lipschitz for each $\mathbf{y}\in\mathcal{Y}$, we have $\|\mathbf{g}\|\leq L$ for all $\mathbf{g}\in\partial_{\mathbf{x}}f(\mathbf{x},\mathbf{y})$ and each pair of $\mathbf{x}$ and $\mathbf{y}\in\mathcal{Y}$. Applying the same argument used for proving Lemma C.2 yields
@@
Applying the same argument used for proving Lemma C.2 yields
@@
Proof. Applying the same argument used for proving Lemma C.3 but with the two inequalities in Lemma D.2, we can obtain the inequalities for Algorithm 1 and 2. $\Box$
@@
We first prove the results for Algorithm 1. Applying the same argument used for proving Theorem 4.6 but with the first inequalities in Lemmas D.1 and D.3, we have
@@
This implies that the number of gradient evaluations required by Algorithm 1 to return an $\epsilon$-stationary point is
@@
We then prove the results for Algorithm 2. Applying the same argument used for analyzing Algorithm 1 but with the second inequalities in Lemmas D.1 and D.3 and using $B=\lfloor\sqrt{\frac{1}{\mu\eta_{\mathbf{x}}}}\rfloor+1$ and the definition of $\eta_{\mathbf{x}}$, we have
@@
This implies that the number of stochastic gradient evaluations required by Algorithm 2 to return an $\epsilon$-stationary point is
@@ Appendix E Nonsmooth and Nonconvex-Concave Setting
Proof. Applying the same argument used for proving Lemma C.3 but with the two inequalities in Lemma E.1, we can obtain the inequalities for Algorithm 1 and 2. $\Box$
@@
We first prove the results for Algorithm 1. Applying the same argument used for proving Theorem 4.6 but with the first inequalities in Lemmas D.1 and E.2, we have
@@
This implies that the number of gradient evaluations required by Algorithm 1 to return an $\epsilon$-stationary point is
@@
We now prove the results for Algorithm 2. Applying the same argument used for analyzing Algorithm 1 but with the second inequalities in Lemmas D.1 and E.2, we have
@@
This implies that the number of stochastic gradient evaluations required by Algorithm 2 to return an $\epsilon$-stationary point is